A point charge q is located at distances a and b from two perpendicular conducting half-planes, both at zero potential. Calculate the force acting on the charge q.
Answer #1

ci by Image metheplan 7- 29 F I b Electrostatic force F = kg,02 az -kq2 - k92 462 492 k q2 9(2²+62 - fret = to - 1 ² + 4² . kIf you have any doubt then mention in comment.